HANDY SEW COMPANY Managerial Finance Concept:                                Managing inventory by minimi

HANDY SEW COMPANY

Managerial Finance Concept:                                Managing inventory by minimizing inventory cost

Financial Analysis Technique:                                Economic order Quantity

Decision Context:                                                      A manufacturer of hand-held sewing machine wants to determine the optimum inventory and safety stock levels for precision parts.

Andy Soo, owner and general manager of Handy Sew Company, a start-up business, was preparing the inventory plan for his company for the next five years. The company was set up to manufacture, assemble and sell hand-held sewing machines. Critical parts of the product were precision components that run the machine under the control of a microchip. Soo believed that the economic order quantity (EOQ) model would be useful in managing the purchasing and inventory of this component.

The Business and Product

Andy Soo, who had a plastics manufacturing business in Manila, saw a business opportunity while visiting Hong Kong. A local manufacturer of a hand-held sewing machine in the colony, Single Machines Company, invited him to run a franchise in the Philippines to assemble and sell its product. After a brief negotiation, the two parties struck an agreement allowing Soo to manufacture the hard-plastic body of the sewing machine and to purchase all precision components from a distributor in Manila. Final assembly of the body and precision parts could be done with a screwdriver.

Soo found the offer an attractive one. He planned to retool a part of his plant to manufacture the hard-plastic body of the product. The technology needed was familiar to him and his workers. His problem was how to market the product and how to control costs.

He believed that there was a growing market for convenient equipment such as a hand-held sewing machine that could sew end mend clothes, bedsheets, curtains and other similar light materials. Travelers would find it useful and could easily pack it in suitcases. It could also be useful at home for mending work. The machine was easy to set up, battery operated, and simple to operate.

The set of precision parts of the machine was a reduced-scale version of that found in standard sewing machines. They could perform the same functions except for limitations in handling heavy textile materials and ill sewing more intricate patterns. These components accounted for more than 80 percent of the total product cost. Their patents were held by Single Machines in Hong Kong, and they were manufactured at a highly automated plant that shipped to a worldwide distribution network.

Due to the importance of these precision parts in his company’s a profitability, Soo wanted to use the economic order quantity model for his purchasing and inventory planning.

Plans for Precision Parts Inventory

The hand-held sewing machine was only one of the many plastic products that Soo was selling. He wanted to start slow while exploring the market. He conducted a study or the likely market, which he believed was the two-income family oriented to travelling for business and pleasure. He also estimated the ordering end carrying costs of the precision parts for the next five years.

He forecast annual requirements for precision parts to increase from 8,000 sets on the first years to 16,000 sets in the fifth year. Soo expected ordering costs to decline from P6,200 per order by the fifth year. The somewhat high ordering cost was due to the requirements by Single Machines for inspection and audit of Handy Sew to ensure that earlier orders went into bona fide production of the sewing machines covered by the agreement.

Soo expected the carrying costs to increase from 6 percent of the average inventory value in the first year to 14 percent in the first year. The projected increases reflected the increasing costs of controlling warehousing losses in high volume operations. The microchips and miniaturized sewing machine parts were sensitive to handling and pilferage losses. Table 1 shows the forecast volumes, ordering costs and carrying costs for the next 5 years.

Table 1. Forecast of Volume Carrying Cost and Ordering Cost for Precision Parts (5 years)

Year

1

2

3

4

6

Volume (sets)

Ordering costs (per order)

Carrying cost (% of average inventory value)

8,000

P6,200

6

9,000

P5,000

8

11,000

P3,800

10

13,000

P2,500

12

16,000

P800

11

The purchase price of one precision parts set was set under the agreement at P3,250 in the first year. A price escalation of 10 percent per year was stipulated in the agreement. The distributor of Single Machines in Manila promised a lead time of 10 days but warned that, due to inspection requirements, it could be as short five days or as long as 14 days.

Soo thought that he could set up a safety stock to meet the problem of highly variable purchasing lead times. Looking at the information he had put together, Soo began to think that the problem might be too complex for him to handle. Before considering the EOQ model, his plan was to purchase the precision parts twice a year and to keep sufficient storage spare to handle the stocks. He wanted to compare the inventory costs under the EOQ model and those under the twice-a-year purchase plan. If the difference was not substantial, he would drop the EOQ model.

Guide Questions

d) If the company purchases precision parts twice a year:

What the average level of inventory of precision parts?

Calculate the total annual inventory costs
Answer #1

d- Average level of inventory in year 1= 8000/2*6%=240, Carrying cost= 240*3250=780.000

Average level of inventory in year 2= 9000/2*8%=360, Carrying cost= 360*3250*1.1=360*3575=1,287.000

Average level of inventory in year 3= 11000/2*10%=550, Carrying cost= 550*3575*1.1=550*3933=2,162,875

Average level of inventory in year 4= 13000/2*12%=780, Carrying cost= 780*3933*1.1=780*4326=3,374,514

Average level of inventory in year 5= 16000/2*11%=880, Carrying cost= 880*4326*1.1=880*4759=4,187,568

Total Annual inventory Cost in Year 1=8000*3250+2*6200+780,000= 26,792,400

Total Annual inventory Cost in Year 2=9000*3575+2*5000+1,287,000=33,472,000

Total Annual inventory Cost in Year 3=11000*3933+2*3800+2,162,875=45,433,475

Total Annual inventory Cost in Year 4=13000*4326+2*2500+3,374,514=59,617,514

Total Annual inventory Cost in Year 5=16000*4759+2*800+4,187,568=80,333,168